# The Quantum Esscher Transform

### v1

@@ 2.2.3 Connection to quantum imaginary time evolution
Next, we discuss how to implement the quantum Esscher Transform on quantum computers using modern techniques based on block-encodings (BE) and the quantum singular value transformation (QSVT). Before doing so we collate the relevant tools and techniques of the framework in the next section.
3 Overview on block-encodings and quantum singular value transformations
The technique of quantum signal processing [LYC16] and its lifting, via ‘qubitization’, to quantum singular value transformation (QSVT) [LC19, GSLW19] provide a concise way to formulate quantum algorithms, particularly for linear algebraic tasks. This framework has provided more efficient implementations of several existing quantum algorithms, such as Hamiltonian simulation [LC17, LC19], amplitude amplification and estimation [GSLW19, RF23] and quantum linear systems solving [GSLW19], and even led to the discovery of new algorithms. For our purposes, we do not actually need the full generality of QSVT. As our matrices of interest are Hermitian and thus admit spectral decompositions, a relaxed version of QSVT—quantum eigenvalue transformation (QET)—suffices. We direct readers interested in learning more about QSVT to [GSLW19, MRTC21, ${}^{+}$23].
Definition 3.1 (Block-Encoding).
Remark 3.2.
[GSLW19] provides a construction of exact block-encodings for density operators, assuming access to oracles which prepare the purifications of the density operators:
Definition 3.3 (Purified quantum query-access).
[…]
Proposition 3.4 (Block-encoding of density operators – Lemma 45, [GSLW19]).
For general matrices which need not be density operators, [CGJ18, GSLW19] also showed how to implement their block-encodings efficiently, assuming the existence of quantum random access memory (QRAM) [GLM08]. Given block-encodings of operators $A_{i}$, we can construct block-encodings of their linear combinations and products. For linear combinations, we make use of an auxiliary tool known as a ‘state preparation pair’. Recall that $\|\cdot\|_{1}$ is the $l_{1}$/Manhattan norm.
Definition 3.5 (State Preparation Pair).
[…]
Proposition 3.6 (Linear combination of block-encoded matrices – Lemma 52, [GSLW19]).
Let
  1. i.
  2. ii.
    $A=\sum_{j=0}^{m-1}y_{j}A_{j}$ for $y:=(y_{0},\dots,y_{m-1})\in\mathbb{C}^{m}$,
  3. iii.
where
[…]
is a $(n+a+b)$-qubit unitary.
In Proposition 3.6, the subnormalization factors of the $A_{j}$’s are to be the same. Later on, we will need a slight generalization of the above result whereby this requirement is dropped.
Proposition 3.7 (Generalized linear combination of block-encoded matrices).
Let
  1. i.
  2. ii.
    $A=\sum_{j=0}^{m-1}y_{j}A_{j}$ for $y:=(y_{0},\dots,y_{m-1})\in\mathbb{C}^{m}$,
  3. iii.
where
[…]
is a $(n+a+b)$-qubit unitary.
Proof.
The following is adapted from the proof of Lemma 52, [GSLW19]. By definition of state-preparation pairs (see Definition 3.5), $P_{L}\ket{0^{b}}=\sum_{j=0}^{2^{b}-1}c_{j}\ket{j}$ and $P_{R}\ket{0^{b}}=\sum_{j=0}^{2^{b}-1}d_{j}\ket{j}$ such that $\sum_{j=0}^{m-1}|\alpha_{j}y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}$. First we evaluate the block extraction of $\widetilde{W}$. We have
$\quad\;(\bra{0^{b+a}}\otimes I_{n})\widetilde{W}(\ket{0^{b+a}}\otimes I_{n})$
[…]
$\leq\varepsilon_{\text{SP}}+\sum_{j=0}^{m-1}|y_{j}|\varepsilon_{\text{BE}}$
$\leq\varepsilon_{\text{SP}}+\frac{\beta}{\inf_{j}\alpha_{j}}\varepsilon_{\text{BE}}.$
Remark 3.8.
In the special case where the block-encodings of the $A_{j}$’s have the same subnormalization factors, i.e., $\alpha_{j}=\alpha$ for all $j$, we recover Proposition 3.6 from Proposition 3.7 . To see this, observe that if $(P_{L},P_{R})$ is a $(\beta,b,\varepsilon_{\text{SP}})$-state-preparation-pair for $\alpha\odot y$, then $\sum_{j}|\alpha_{j}y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}\implies\sum_{j}|\alpha y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}\implies\sum_{j}|y_{j}-\frac{\beta}{\alpha}c_{j}^{*}d_{j}|\leq\frac{\varepsilon_{\text{SP}}}{\alpha}$, thus implying $(P_{L},P_{R})$ is a $(\frac{\beta}{\alpha},b,\frac{\varepsilon_{\text{SP}}}{\alpha})$-state-preparation-pair for $y$. According to Proposition 3.6, $\widetilde{W}$ is then a $(\alpha\cdot\frac{\beta}{\alpha},\;a+b,\;\alpha\cdot\frac{\varepsilon_{\text{SP}}}{\alpha}+\frac{\beta}{\alpha}\varepsilon_{\text{BE}})$-BE of $A$. This is in agreement with Proposition 3.7.
Theorem 3.9 (Polynomial Eigenvalue Transformation – Theorem 56, [GSLW19]).
Proposition 3.10 (Bounded Polynomial Approximation – Corollary 66, [GSLW19]).
$\|f(x)-P(x)\|_{[x_{0}-r,x_{0}+r]}$ $\leq\varepsilon$ (3.1)
$\|P(x)\|_{[-1,1]}$ $\leq\varepsilon+\|f(x)\|_{[x_{0}-r-\delta/2,x_{0}+r+\delta/2]}\leq\varepsilon+B$ (3.2)
$\|P(x)\|_{[-1,1]\setminus[x_{0}-r-\delta/2,x_{0}+r+\delta/2]}$ $\leq\varepsilon.$ (3.3)
Theorem 3.9 and Proposition 3.10 are to be used in conjunction to produce block-encodings of general functions of Hermitian matrices. In doing so, we first note that Theorem 3.9 produces an encoding of $P(A/\alpha)$, not $P(A)$. Thus, with a polynomial approximation of $f$, say $P(x)\approx f(x)$, it is generally not true that $P(A/\alpha)\approx f(A)$. What we need is a polynomial approximation not of $f$, but of a (horizontally) scaled version of $f$, $f^{\prime}(x):=f(\alpha x)$, so that $P(x)\approx f^{\prime}(x)\implies P(A/\alpha)\approx f^{\prime}(A/\alpha)=f(A)$. Second, we also have to take into account the polynomial approximation error incurred in producing the final desired block encoding $f(A)$. We take care of these matters in Corollary 3.11, which, given the block-encoding of an arbitrary Hermitian matrix $A$, produces a block-encoding of $f(A)$, where $f$ is a generic real-valued function.
Corollary 3.11 (Block-encoding functions of general Hermitian matrices).
Given
  1. i.
  2. ii.
    $f:I\longrightarrow\mathbb{R}$, a smooth function on an open interval $I$ containing $[\lambda_{\min},\lambda_{\max}]$. Assume the function $x\mapsto f(\alpha x)$ satisfies the conditions in Proposition 3.10 with $[\lambda_{\min}/\alpha,\lambda_{\max}/\alpha]\subseteq[x_{0}-r,x_{0}+r]$ and series-of-coefficients bound $B$.
  3. iii.
Proof.
By Proposition 3.10, given polynomial approximation error tolerance $\varepsilon_{\text{poly}}$ there exists a polynomial $Q\in\mathbb{C}[x]$ of degree $\mathcal{O}\left(\frac{1}{\delta}\log\left(\frac{B}{\varepsilon_{\text{poly}}}\right)\right)$ which $\varepsilon_{\text{poly}}$-approximates $f\circ t_{\alpha}^{-1}$ on $[x_{0}-r,x_{0}+r]$ and is bounded above by $1+B$ on $[-1,1]$. Since $\|A/\alpha\|\in[\lambda_{\min}/\alpha,\lambda_{\max}/\alpha]\subseteq[x_{0}-r,x_{0}+r]$, we have
In order to apply Theorem 3.9, our polynomial has to be real and upper-bounded by $1/2$ on $[-1,1]$. Observe that for any complex-valued function $F$ and domain $S$,
Since $f$ itself is real-valued, $\operatorname{Re}Q\in\mathbb{R}[x]$ is qualified to assume the role of $P$ in Proposition 3.10. That is, the real polynomial $\operatorname{Re}Q$ also $\varepsilon_{\text{poly}}$-approximates $f\circ t_{\alpha}^{-1}$ on $[x_{0}-r,x_{0}+r]$ and is bounded above by $1+B$ on $[-1,1]$. Thus, letting $P\leftarrow\frac{\operatorname{Re}Q}{2(1+B)}$ in Theorem 3.9 we obtain $\tilde{U}$, a $(1,a+2,4d\sqrt{\varepsilon/\alpha})$-encoding of $\frac{\operatorname{Re}Q}{2(1+B)}(A/\alpha)$, where $d=\mathcal{O}\left(\frac{1}{\delta}\log\left(\frac{B}{\varepsilon_{\text{poly}}}\right)\right)$. Putting these together and noting that $f\circ t_{\alpha}^{-1}(\frac{A}{\alpha})=f(A)$, we have
$\left\|\frac{f(A)}{2(1+B)}-(\bra{0^{a+2}}\otimes I)\tilde{U}(\ket{0^{a+2}}\otimes I)\right\|$
$\qquad\leq\left\|\frac{f\circ t_{\alpha}^{-1}(\frac{A}{\alpha})}{2(1+B)}-\frac{\operatorname{Re}Q(\frac{A}{\alpha})}{2(1+B)}\right\|+\left\|\frac{\operatorname{Re}Q(\frac{A}{\alpha})}{2(1+B)}-(\bra{0^{a+2}}\otimes I)\tilde{U}(\ket{0^{a+2}}\otimes I)\right\|$
$\qquad\leq\frac{\varepsilon_{\text{poly}}}{2(1+B)}+4d\sqrt{\varepsilon/\alpha}.$
4 Implementation on quantum computers
@@
Reference [GSLW19] demonstrates how to construct block-encodings for density operators $\rho$ within the purified quantum query-access model (see Definition 3.3 and Proposition 3.4 above). For the Hermitian operators $H_{i}$ which are generally not density operators, their block-encodings can be constructed efficiently for many physical Hamiltonians, or if the $H_{i}$’s are stored in sparse data structures or KP trees. Along the way we shall also need as an auxiliary tool ‘state-preparation pairs’ (see Definition 3.5), to prepare linear combinations of the Hamiltonians. We assume immediate access to these, as we do for block-encodings. For the construction of state-preparation pairs, one can refer to [vAG18].
4.1 Technical lemmas
@@
Lemma 4.1 (Block-encoding of $\log\rho$).
@@
First we construct a polynomial approximation of $\log x$. More specifically, we check that the function $\log x$ satisfies the conditions of Proposition 3.10, with the appropriate $x_{0},r,\delta$ and $B$. Corollary 3.11 then gives us the desired block-encoding.
The following derivation is based on the proof of Corollary 67, [GSLW19] and Lemma 11, [GL19]. Negative power functions $x^{-c}$ share with $\log x$ the common property of going to infinity as $x$ approaches 00, thus the Taylor expansions of these functions are performed about $x=1$. Choose $x_{0}=1$, $r=1-\frac{1}{\kappa}$ and $\delta=\frac{1}{2\kappa}$. The Taylor series of $\log x$ about $x=1$ is $\log x=\sum_{k=1}^{\infty}\frac{(-1)^{k+1}}{k}(x-1)^{k}$. With $a_{k}=\frac{(-1)^{k+1}}{k}$, the series-of-coefficients bound $B$ in Proposition 3.10 is
$\sum_{k=1}^{\infty}(r+\delta)^{k}|a_{k}|=\sum_{k=1}^{\infty}\frac{(1-1/2\kappa)^{k}}{k}=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\left(\frac{1}{2\kappa}-1\right)^{k}=-\log\frac{1}{2\kappa}=\log 2\kappa=:B.$
Corollary 3.11 gives us the unitary $U_{\log\rho}$, which is a $\left(2(1+\log 2\kappa),\;a+2,\;\varepsilon_{\text{poly}}\right)$-encoding of $\log\rho$, which can be constructed using $\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{poly}}}\right)\right)$ queries to $U_{\rho}$. ∎
@@
Lemma 4.2 (Block-encoding of $e^{H}$).
@@
By Corollary 64, [GSLW19], there exists $P\in\mathbb{R}[x]$ of degree $t=\mathcal{O}\left(\sqrt{\max(\alpha,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)$ such that $\|\frac{e^{\alpha x}}{e^{\alpha}}-P(x)\|_{[-1,1]}\leq\varepsilon_{\text{poly}}$. Furthermore $\|P(x)\|\leq\|\frac{e^{\alpha x}}{e^{\alpha}}-P(x)\|_{[-1,1]}+\|\frac{e^{\alpha x}}{e^{\alpha}}\|_{[-1,1]}\leq 1+B$, where $B=1$. Applying Corollary 3.11 with $f(x)=\frac{e^{x}}{e^{\alpha}}$ gives a $\left(4,\;a+2,\;\varepsilon_{\text{poly}}+16t\sqrt{\varepsilon/\alpha}\right)$-encoding of $e^{H}/e^{\alpha}$, making $t$ queries to $U_{H}$. ∎
4.2 Algorithm
We now provide the algorithm implementing the quantum Esscher transform, see Algorithm 1. We specify the constraints on the inputs and the guarantees on the output in the algorithm itself. A step-by-step analysis of Algorithm 1 is provided below in detail, whereafter the overall (query) complexity is stated. We summarize these information in Theorem 4.3.
Theorem 4.3.
queries to $U_{\rho}$ and
$\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$
queries to each $U_{j}$.
@@ 4.2 Algorithm
8:Construct $U_{\log\rho}$, a $(2(1+\log 2\kappa),\;n+n_{\rho}+2,\;\varepsilon_{\text{BE}})$-BE of $\log\rho$. This makes $t=\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{BE}}}\right)\right)$ queries to $U_{\rho}$, see Lemma 4.1.
@@
13:Using $(P_{L},P_{R})$, combine $U_{\log\rho}$ and $U_{j}$, $j\in[d]$ to give $U_{H}$, a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H:=\sum_{i}\theta_{i}H_{i}+\log\rho$. This makes 1 query to $(P_{L},P_{R})$ and 1 query to $U_{\log\rho}$ and each $U_{j}$, see Proposition 3.7.
14:Construct $U_{\sigma}$, a $(1,\;\max\{a,n+n_{\rho}\}+4+\lceil\log d\rceil,\;\varepsilon)$-BE of $\sigma:=e^{H}/\mathcal{N}$. Makes $t=\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$ queries to $U_{H}$, see Lemma 4.2.
@@
Proof of Theorem 4.3.
@@
Step 1. From Proposition 3.4 we construct $U_{\rho}=\widetilde{O_{\rho}}:=(O_{\rho}^{\dagger}\otimes I_{n})(I_{n+n_{\rho}}\otimes\text{SWAP}_{n})(O_{\rho}\otimes I_{n})$, a $(1,n+n_{\rho},0)$-BE of $\rho$. This makes $\mathcal{O}(1)$ queries to $O_{\rho}$.
@@
Step 2. This step entails a polynomial approximation to the logarithm function on the interval $[\frac{1}{\kappa},1]$. Denote by $\varepsilon_{\text{poly}}$ the approximation error tolerance. Choose $\varepsilon_{\text{poly}}\leq\varepsilon_{\text{BE}}$. Lemma 4.1 gives $U_{\log\rho}$, a $(2(1+\log 2\kappa),\;n+n_{\rho}+2,\;\varepsilon_{\text{BE}})$-BE of $\log\rho$. The construction of $U_{\log\rho}$ makes $t=\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{BE}}}\right)\right)$ queries to $U_{\rho}$, where $t$ is the degree of the approximating polynomial (see Proposition 3.10/Corollary 3.11).
@@
Step 3. Construct a $(\beta,b,\varepsilon_{\text{SP}})$-state-preparation-pair $(P_{L},P_{R})$ for $\alpha\odot\theta\in\mathbb{R}^{d+1}$, where $\alpha=(1^{d},2(1+\log 2\kappa))$ and $\theta=(\theta_{1},\dots,\theta_{d},1)$ (see Proposition 3.7). Choose $\beta=\|\alpha\odot\theta\|_{1}=\|\theta\|_{1}+2(1+\log 2\kappa)$. $b$ has to be such that $d+1\leq 2^{b}$, so choose $b=\lceil\log d\rceil$. Finally, choose $\varepsilon_{\text{SP}}\leq\beta\varepsilon_{\text{BE}}$. The construction of $(P_{L},P_{R})$ can be achieved using $\mathcal{O}(d)$ elementary gates [${}^{+}$15].
@@
Step 4. Now we make use of our access to the state-preparation-pair $(P_{L},P_{R})$. To form linear combinations of block-encodings, the number of ancilla qubits required for each constituent block-encoding should be the same, see Proposition 3.6/3.7. Remark 3.2 shows that we can always equalize this number of ancilla qubits by padding with additional ancillas. The equalized number of ancillas is $\max\{a,n+n_{\rho}+2\}\leq\max\{a,n+n_{\rho}\}+2$. We could also take $a+n+n_{\rho}+2$, but we want to minimize the number of ancilla qubits. From Proposition 3.7 we get $U_{H}$, a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H:=\sum_{i}\theta_{i}H_{i}+\log\rho$, making 1 query to $(P_{L},P_{R})$ and 1 query to $U_{\log\rho}$ and each $U_{j}$.
@@
Step 5. Finally, we construct a block-encoding for $e^{H}/\mathcal{N}$. At this stage, we have a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H$. Lemma 4.2 gives a $(1,\;\max\{a,n+n_{\rho}\}+\lceil\log d\rceil+4,\;\varepsilon_{\text{poly}}/4+4t\sqrt{2\varepsilon_{\text{BE}}})$-BE of $\sigma=e^{H}/{4e^{\beta}}$ (thus $\mathcal{N}=4e^{\beta}$), where $t=\mathcal{O}\left(\sqrt{\max(\beta,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)$. It remains to make judicious choices for $\varepsilon_{\text{poly}}$ (note that the $\varepsilon_{\text{poly}}$ at this step need not be the same as the one in Step 2) and $\varepsilon_{\text{BE}}$ in order to ensure the overall block-encoding error is less than $\varepsilon$, i.e.
$\frac{\varepsilon_{\text{poly}}}{4}+4t\sqrt{2\varepsilon_{\text{BE}}}\leq\varepsilon.$ (4.1)
@@
These choices ensure Equation 4.1 is satisfied. Note that $\lim_{x\rightarrow 0}\frac{x}{\log\frac{1}{x}}=0$, so $\varepsilon_{\text{BE}}\rightarrow 0$ as $\varepsilon\rightarrow 0$. The degree of the approximating polynomial, and thus the number of queries to $U_{H}$ required, is $t=\mathcal{O}\left(\sqrt{\max(\beta,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)=\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$. Recall that constructing $U_{H}$ itself makes 1 query to $U_{\log\rho}$ and each $U_{j}$. Lastly, observe that $\|e^{H}\|\leq e^{\|H\|}\leq e^{\sum_{i}|\theta_{i}|+\log\kappa}\leq e^{\beta}<\mathcal{N}$, so $\mathcal{N}$ is a valid subnormalization factor.
@@
$\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)\cdot\mathcal{O}\left(\kappa\log\left(\log\kappa\cdot\frac{1}{\varepsilon^{2}}\cdot\log^{2}\frac{1}{\varepsilon}\right)\right)\subseteq\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon}\right)\log\left(\frac{1}{\varepsilon}\right)\right)\subseteq\widetilde{\mathcal{O}}\left(\kappa\log^{2}\left(\frac{1}{\varepsilon}\right)\right)$

### v2

@@ II.2 Quantum version
Next, we discuss how to implement the quantum Esscher Transform on quantum computers using modern techniques based on block-encodings (BE) and the quantum singular value transformation (QSVT). The relevant tools and techniques of the framework are collated in the appendix.
III Implementation on quantum computers
@@
Reference [15] demonstrates how to construct block-encodings for density operators $\rho$ within the purified quantum query-access model (see Definition 4 and Proposition 12 below). For the Hermitian operators $H_{i}$ which are generally not density operators, their block-encodings can be constructed efficiently for many physical Hamiltonians, or if the $H_{i}$’s are stored in sparse data structures or KP trees. Along the way we shall also need as an auxiliary tool ‘state-preparation pairs’ (see Definition 5), to prepare linear combinations of the Hamiltonians. We assume immediate access to these, as we do for block-encodings. For the construction of state-preparation pairs, one can refer to [36].
III.1 Technical lemmas
@@
Lemma 3 (Block-encoding of $\log\rho$).
@@
First we construct a polynomial approximation of $\log x$. More specifically, we check that the function $\log x$ satisfies the conditions of Proposition 16, with the appropriate $x_{0},r,\delta$ and $B$. Corollary 17 then gives us the desired block-encoding.
The following derivation is based on the proof of Corollary 67, [15] and Lemma 11, [11]. Negative power functions $x^{-c}$ share with $\log x$ the common property of going to infinity as $x$ approaches 00, thus the Taylor expansions of these functions are performed about $x=1$. Choose $x_{0}=1$, $r=1-\frac{1}{\kappa}$ and $\delta=\frac{1}{2\kappa}$. The Taylor series of $\log x$ about $x=1$ is $\log x=\sum_{k=1}^{\infty}\frac{(-1)^{k+1}}{k}(x-1)^{k}$. With $a_{k}=\frac{(-1)^{k+1}}{k}$, the series-of-coefficients bound $B$ in Proposition 16 is
$\sum_{k=1}^{\infty}(r+\delta)^{k}|a_{k}|$ $=$ $\sum_{k=1}^{\infty}\frac{(1-\frac{1}{2\kappa})^{k}}{k}=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\left(\frac{1}{2\kappa}-1\right)^{k}$ (III.1)
$=$ $-\log\frac{1}{2\kappa}=\log 2\kappa=:B.$
Corollary 17 gives us the unitary $U_{\log\rho}$, which is a $\left(2(1+\log 2\kappa),\;a+2,\;\varepsilon_{\text{poly}}\right)$-encoding of $\log\rho$, which can be constructed using $\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{poly}}}\right)\right)$ queries to $U_{\rho}$. ∎
@@
Lemma 4 (Block-encoding of $e^{H}$).
@@
By Corollary 64, [15], there exists $P\in\mathbb{R}[x]$ of degree $t=\mathcal{O}\left(\sqrt{\max(\alpha,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)$ such that $\|\frac{e^{\alpha x}}{e^{\alpha}}-P(x)\|_{[-1,1]}\leq\varepsilon_{\text{poly}}$. Furthermore $\|P(x)\|\leq\|\frac{e^{\alpha x}}{e^{\alpha}}-P(x)\|_{[-1,1]}+\|\frac{e^{\alpha x}}{e^{\alpha}}\|_{[-1,1]}\leq 1+B$, where $B=1$. Applying Corollary 17 with $f(x)=\frac{e^{x}}{e^{\alpha}}$ gives a $\left(4,\;a+2,\;\varepsilon_{\text{poly}}+16t\sqrt{\varepsilon/\alpha}\right)$-encoding of $e^{H}/e^{\alpha}$, making $t$ queries to $U_{H}$. ∎
@@
8:Construct $U_{\log\rho}$, a $(2(1+\log 2\kappa),\;n+n_{\rho}+2,\;\varepsilon_{\text{BE}})$-BE of $\log\rho$. This makes $t=\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{BE}}}\right)\right)$ queries to $U_{\rho}$, see Lemma 3.
@@
13:Using $(P_{L},P_{R})$, combine $U_{\log\rho}$ and $U_{j}$, $j\in[d]$ to give $U_{H}$, a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H:=\sum_{i}\theta_{i}H_{i}+\log\rho$. This makes 1 query to $(P_{L},P_{R})$ and 1 query to $U_{\log\rho}$ and each $U_{j}$, see Proposition 14.
14:Construct $U_{\sigma}$, a $(1,\;\max\{a,n+n_{\rho}\}+4+\lceil\log d\rceil,\;\varepsilon)$-BE of $\sigma:=e^{H}/\mathcal{N}$. Makes $t=\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$ queries to $U_{H}$, see Lemma 4.
@@
III.2 Algorithm
We now provide the algorithm implementing the quantum Esscher Transform, see Algorithm 1. We specify the constraints on the inputs and the guarantees on the output in the algorithm itself. A step-by-step analysis of Algorithm 1 is provided below in detail, whereafter the overall (query) complexity is stated. We summarize these information in Theorem 5.
Theorem 5.
queries to $U_{\rho}$ and
$\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$
queries to each $U_{j}$.
Proof of Theorem 5.
@@ III.2 Algorithm
From Proposition 12 we construct $U_{\rho}=\widetilde{O_{\rho}}:=(O_{\rho}^{\dagger}\otimes I_{n})(I_{n+n_{\rho}}\otimes\text{SWAP}_{n})(O_{\rho}\otimes I_{n})$, a $(1,n+n_{\rho},0)$-BE of $\rho$. This makes $\mathcal{O}(1)$ queries to $O_{\rho}$.
Step 2
@@
This step entails a polynomial approximation to the logarithm function on the interval $[\frac{1}{\kappa},1]$. Denote by $\varepsilon_{\text{poly}}$ the approximation error tolerance. Choose $\varepsilon_{\text{poly}}\leq\varepsilon_{\text{BE}}$. Lemma 3 gives $U_{\log\rho}$, a $(2(1+\log 2\kappa),\;n+n_{\rho}+2,\;\varepsilon_{\text{BE}})$-BE of $\log\rho$. The construction of $U_{\log\rho}$ makes $t=\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon_{\text{BE}}}\right)\right)$ queries to $U_{\rho}$, where $t$ is the degree of the approximating polynomial (see Proposition 16/Corollary 17).
Step 3
@@
Construct a $(\beta,b,\varepsilon_{\text{SP}})$-state-preparation-pair $(P_{L},P_{R})$ for $\alpha\odot\theta\in\mathbb{R}^{d+1}$, where $\alpha=(1^{d},2(1+\log 2\kappa))$ and $\theta=(\theta_{1},\dots,\theta_{d},1)$ (see Proposition 14). Choose $\beta=\|\alpha\odot\theta\|_{1}=\|\theta\|_{1}+2(1+\log 2\kappa)$. $b$ has to be such that $d+1\leq 2^{b}$, so choose $b=\lceil\log d\rceil$. Finally, choose $\varepsilon_{\text{SP}}\leq\beta\varepsilon_{\text{BE}}$. The construction of $(P_{L},P_{R})$ can be achieved using $\mathcal{O}(d)$ elementary gates [2].
Step 4
@@
Now we make use of our access to the state-preparation-pair $(P_{L},P_{R})$. To form linear combinations of block-encodings, the number of ancilla qubits required for each constituent block-encoding should be the same, see Proposition 13/14. Remark 3 shows that we can always equalize this number of ancilla qubits by padding with additional ancillas. The equalized number of ancillas is $\max\{a,n+n_{\rho}+2\}\leq\max\{a,n+n_{\rho}\}+2$. We could also take $a+n+n_{\rho}+2$, but we want to minimize the number of ancilla qubits. From Proposition 14 we get $U_{H}$, a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H:=\sum_{i}\theta_{i}H_{i}+\log\rho$, making 1 query to $(P_{L},P_{R})$ and 1 query to $U_{\log\rho}$ and each $U_{j}$.
Step 5
@@
Finally, we construct a block-encoding for $e^{H}/\mathcal{N}$. At this stage, we have a $(\beta,\;\max\{a,n+n_{\rho}\}+2+\lceil\log d\rceil,\;2\beta\varepsilon_{\text{BE}})$-BE of $H$. Lemma 4 gives a $(1,\;\max\{a,n+n_{\rho}\}+\lceil\log d\rceil+4,\;\varepsilon_{\text{poly}}/4+4t\sqrt{2\varepsilon_{\text{BE}}})$-BE of $\sigma=e^{H}/{4e^{\beta}}$ (thus $\mathcal{N}=4e^{\beta}$), where $t=\mathcal{O}\left(\sqrt{\max(\beta,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)$. It remains to make judicious choices for $\varepsilon_{\text{poly}}$ (note that the $\varepsilon_{\text{poly}}$ at this step need not be the same as the one in Step 2) and $\varepsilon_{\text{BE}}$ in order to ensure the overall block-encoding error is less than $\varepsilon$, i.e.
$\frac{\varepsilon_{\text{poly}}}{4}+4t\sqrt{2\varepsilon_{\text{BE}}}\leq\varepsilon.$ (III.2)
@@
These choices ensure Equation III.2 is satisfied. Note that $\lim_{x\rightarrow 0}\frac{x}{\log\frac{1}{x}}=0$, so $\varepsilon_{\text{BE}}\rightarrow 0$ as $\varepsilon\rightarrow 0$. The degree of the approximating polynomial, and thus the number of queries to $U_{H}$ required, is $t=\mathcal{O}\left(\sqrt{\max(\beta,\log\frac{1}{\varepsilon_{\text{poly}}})\log\frac{1}{\varepsilon_{\text{poly}}}}\right)=\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$. Recall that constructing $U_{H}$ itself makes 1 query to $U_{\log\rho}$ and each $U_{j}$. Lastly, observe that $\|e^{H}\|\leq e^{\|H\|}\leq e^{\sum_{i}|\theta_{i}|+\log\kappa}\leq e^{\beta}<\mathcal{N}$, so $\mathcal{N}$ is a valid subnormalization factor.
@@
$\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)\cdot\mathcal{O}\left(\kappa\log\left(\log\kappa\cdot\frac{1}{\varepsilon^{2}}\cdot\log^{2}\frac{1}{\varepsilon}\right)\right)$
$\subseteq\mathcal{O}\left(\kappa\log\left(\frac{\log\kappa}{\varepsilon}\right)\log\left(\frac{1}{\varepsilon}\right)\right)\subseteq\widetilde{\mathcal{O}}\left(\kappa\log^{2}\left(\frac{1}{\varepsilon}\right)\right)$